Exercise 4.9. Let X ~ Poisson(10). (a) Find P(X>7). (b) Find P(X < 13 X > 7).
Answer #1

Exercise .9 Let X poissonCIo) P(x-x) a.) Find Px7) PIx -Px L6) PCX7)= PCX o)+ PCX = 1) + P(x =2) +P(X=3)+ P(x= le) +PLx=5)+PC